Bill is biking across state. It takes 23 hours to complete the 213 mile ride. For the first leg he is pedaling downhill and is averaging the speed of 12 miles an hour. Upon reaching the second leg his speed is cut down to 3 miles an hour. How long did he spend biking downhill?
Answer by lwsshak3(11628)

Bill is biking across state. It takes 23 hours to complete the 213 mile ride. For the first leg he is pedaling downhill and is averaging the speed of 12 miles an hour. Upon reaching the second leg his speed is cut down to 3 miles an hour. How long did he spend biking downhill?
***
let x=time Bill spent biking downhill on the 1st leg of the trip
23-x=time Bill spent on 2nd leg of the trip
distance=speed*travel time
..
12x+3(23-x)=213
12x+69-3x=213
9x=144
x=16
time Bill spent biking downhill=16 hrs